# Reblurring-Guided Single Image Defocus Deblurring: A Learning Framework with Misaligned Training Pairs

### v1

@@ 4.4 Evaluation on DED Dataset
Figure 10: Visual comparison of defocus blur maps. Since defocus blur map cannot be truly captured using a camera, it in this case is estimated from light field data captured by a Lytro camera. The defocus blur map estimated by Ours better reflects spatially variant blur amounts than that predicted by DID-ANet [19].
@@
DID-ANet [19] 29.59 0.860 0.186
Restormer[15] 31.29 0.890 0.121
Ours 31.35 0.893 0.117
@@ 4.5 Generalization Evaluation
This section begins with an evaluation of the performance of models pre-trained on the DPDD and DED datasets, i.e., Restormer from Table 2 and Restormer from Table 3, when applied to haneld images from our SDD dataset. Although defocus blur in the input blurry image is not severe as depicted in Figure 11, both Restormer models trained on DPDD and DED datasets are limited in removing defocus blur, due to the domain gaps between different sensors for capturing images. The limited generalization ability of trained models to new sensors emphasizes the necessity for a rapid data collection and training framework to enhance model adaptability and performance across diverse sensor types.
Furthermore, we follow existing works to adopt RealDOF dataset for evaluating generalization ability. We used the trained model from the DPDD dataset for evaluation. As shown in Fig. 1, the RealDOF dataset also exhibits a certain degree of misalignment. Therefore, the strategy we designed for misaligned datasets also demonstrates effectiveness on the RealDOF dataset. Our evaluation results were calculated between the input images and the ground-truth images after warping, as presented in Table 4. Although trained on the DPDD dataset, Our JDRL framework shows robust generalization to other datasets. The experimental results also indicate that the domain gap between RealDOF and DPDD is smaller compared to the gap between DPDD and SDD. Through our framework, training with loosely aligned data pairs not only reduces data collection costs but also achieves better visual performance, providing an efficient and low-cost deblurring strategy for rapid adaptation to new sensors.
@@
JNB [9] 21.02 0.651 0.590
EBDB [48] 21.37 0.659 0.579
DMENet [49] 21.38 0.658 0.581
DPDNetS [12] 22.66 0.702 0.397
Son et al.[24] 24.35 0.756 0.308
IFAN [14] 25.39 0.794 0.264
MPRNet [51] 24.66 0.771 0.313
Restormer[15] 25.80 0.815 0.249
Loformer[22] 24.72 0.780 0.306
GRLS-B[18] 25.06 0.797 0.261
 Ours 25.95 0.818 0.245
@@ 4.6 Ablation Study
To conduct an efficient ablation study, we employ the UNet as our deblurring model $\mathcal{F}$ to primarily validate the effectiveness of our reblurring-guided learning framework. The effectiveness of the deblurring model involving pseudo defocus blur map for supervision has already been demonstrated in Table 1, where Ours with Eq. (2) outperforms Ours with Eq. (1). Our ablation studies are performed on the SDD dataset, and six variants are designed to analyze their contributions. Variant #1 represents a vanilla UNet trained using $\mathcal{L}_{1}$ loss. Variants #2 and #3 correspond to Ours with Eq. (1) by excluding the bi-directional optical flow deforming process and calibration mask, respectively. Variants #4, #5, and #6 aim to demonstrate the effectiveness of components within our reblurring module. Specifically, they variants of Ours with Eq. (1) by removing the reblurring module, isotropic kernels module $\mathcal{R}_{kpn}$ and weight prediction module $\mathcal{R}_{wpn}$ from, respectively.
@@
#2 w/o Cycle Deformation 24.80/0.758/0.347
#3 w/o Calibration Mask 25.78/0.780/0.309
#4 w/o reblurring module 25.58/0.777/0.311
#5 w/o $\mathcal{R}_{kpn}$ 25.68/0.778/0.311
#6 w/o $\mathcal{R}_{wpn}$ 25.74/0.778/0.332
       Ours (Eq. (1)) 25.94/0.791/0.289
The quantitative results are presented in Table 5. Observations from experiments #1 and #2 demonstrate that the absence of the deformation process significantly diminishes network performance. Moreover, the inclusion of the calibration mask appears to moderately enhance performance, as depicted in experiment #3. Experiment #4 clearly indicates that the lack of reblurring module substantially impacts overall network performance. The results of experiments #5 and #6 validate the advantages of isotropic blur kernels prediction module $\mathcal{R}_{kpn}$ and weight prediction module $\mathcal{R}_{wpn}$ within reblurring module for the reblurring process.

### v2

@@ 4.4 Evaluation on DED Dataset
Figure 12: Visual comparison of defocus blur maps. Since defocus blur map cannot be truly captured using a camera, it in this case is estimated from light field data captured by a Lytro camera. The defocus blur map estimated by Ours better reflects spatially variant blur amounts than that predicted by DID-ANet [19]. Note that our estimated defocus blur maps are with $H\times W\times M$ dimension with $M=35$, and we reduced their dimensionality to $H\times W\times 1$ by adopting principal component analysis.
@@
DID-ANet [19] 29.59 0.860 0.186 72.97 0.193
Restormer[15] 31.29 0.890 0.121 58.74 0.141
$\text{Restormer}^{\dagger}$ 31.37 0.894 0.112 57.26 0.132
@@ 4.5 Generalization Evaluation
Evaluation of Pre-trained models on Different Dataset. This section begins with an evaluation of the performance of models pre-trained on the DPDD [1] and DED datasets[19], i.e., Restormer from Table 3 and Restormer from Table 4, when applied to images from our SDD dataset. Although defocus blur in the input blurry image is not severe as depicted in Figure 13, both Restormer models trained on DPDD and DED datasets are limited in removing defocus blur, due to the domain gaps between different sensors for capturing images. The limited generalization ability of trained models to new sensors emphasizes the necessity for a rapid data collection and training framework to enhance model adaptability and performance across diverse sensor types.
Evaluation of Generalization Ability. Furthermore, we follow existing works to adopt RealDOF dataset for evaluating generalization ability. We used the trained model from the DPDD dataset [1] for evaluation. As shown in Fig. 1, the RealDOF dataset also exhibits a certain degree of misalignment. Therefore, the strategy we designed for misaligned datasets also demonstrates effectiveness on the RealDOF dataset. Our evaluation results were calculated between the input images and the ground-truth images after warping, as presented in Table 5. Although trained on the DPDD dataset, Our learning framework shows robust generalization to other datasets. The experimental results also indicate that the domain gap between RealDOF and DPDD [1] is smaller compared to the gap between DPDD [1] and SDD. Through our framework, training with loosely aligned data pairs not only reduces data collection costs but also achieves better visual performance, providing an efficient and low-cost deblurring strategy for rapid adaptation to new sensors.
Cross-dataset Evaluation. Moreover, we perform cross-dataset evaluation. Within our reblurring-based learning framework, we utilize the UNet architecture to train models on three datasets, i.e., SDD, DPDD [1], and DED [19]. As shown in Table 6, the model trained on DPDD [1] achieves the best performance, followed by our SDD-trained model, while the DED-trained model exhibited the lowest performance. We attribute the observed experimental results primarily to the varying scales and data distribution of the training datasets. Specifically, DPDD [1] comprises 7,000 training pairs, DED [19] contains 1,112 pairs, and our SDD includes 4,830 pairs. We note that the primary contribution of our work lies in providing an approach for effective deployment on target devices by learning a device-specific image defocus deblurring model, while relaxing the strict requirement for perfect alignment in training pairs. Thus, our SDD serves as a testbed for addressing misalignment issues in training pairs and validating the effectiveness of the deblurring models and training strategies, without requiring generalization across different cameras. Consequently, it is acceptable that the model trained on our SDD (captured using a HUAWEI camera) does not achieve the best metrics on RealDOF (captured using a Google Pixel camera).
@@
DPDNetS [1] 22.66 0.702 0.397 73.07 0.213
Son et al.[24] 24.35 0.756 0.308 64.57 0.187
IFAN [2] 25.39 0.794 0.264 43.39 0.162
MPRNet [51] 24.66 0.771 0.313 52.96 0.186
Restormer[15] 25.80 0.815 0.249 40.40 0.153
Loformer[22] 24.72 0.780 0.306 47.99 0.164
$\text{Restormer}^{\dagger}$ 25.98 0.819 0.220 39.69 0.151
Table 6: Evaluation on RealDOF [2] by training UNet models on different datasets.
[…]
DED[19] 21.81 0.672 0.414 79.20 0.221
DPDD[1] 22.28 0.687 0.401 74.27 0.216
SDD 22.03 0.680 0.406 75.41 0.219
@@ 4.6 Ablation Study
To systematically validate the effectiveness of our proposed deblurring framework, we conduct comprehensive ablation studies from two perspectives: (1) the fusion model and (2) the reblurring-guided learning components. All experiments are performed on the SDD dataset using the UNet architecture as the backbone, with quantitative comparisons against variants that systematically remove key modules. These studies aim to isolate the contributions of each component and verify the necessity of our design choices.
4.6.1 Fusion Model Effectiveness Analysis
To validate the superiority of our deformable cross-attention fusion model, we provide experimental comparison against these methods, i.e., DConv [52], LDConv [53], and DAT [54].
Table 7: Comparison of different deformable modules for fusing defocus blur map and blurry image.
[…]
DConv 26.01 0.787 0.303 57.68 0.183
LDConv 25.89 0.774 0.305 58.14 0.184
Ours 26.20 0.796 0.298 55.49 0.180
  • DConv learns per-position offsets within a fixed $k\times k$ grid, adapting locally to geometric variations in single-modality features. While we also employ learned offsets, we apply them in the Transformer attention domain, using query-driven reference points rather than a fixed convolution grid. This allows for more flexible, spatially-global alignment during fusion of blurred images and defocus maps.
  • LDConv extends deformable convolution by enabling arbitrary sampling patterns and linear scaling of filter parameters, offering parameter-efficient but flexible sampling. Unlike LDConv’s convolutional sampling and interpolation, our method leverages Transformer cross-attention to fuse features between modalities using deformable sampling guided by cross-modality queries—providing a more expressive and context-aware fusion mechanism.
  • DAT learns shared offset groups to shift key/value positions from a uniform reference grid, leading to data-dependent, sparse attention. We similarly employ query-conditioned deformable attention, but specifically tailor it to spatial alignment across two modalities (blurry image & defocus map). Our reference points are explicitly optimized to correct misalignment in defocus estimation, rather than general semantic aggregation.
As shown in Table 7, DConv and LDConv exhibit inferior performance compared to Ours, due to their limited capability in capturing long-range correspondence. Overall, deformable attention is likely to outperform deformable convolution, especially in scenarios with significant spatial misalignment, owing to its broader perception field. In Table 7, the results of DAT are omitted, because of its excessive GPU memory consumption, such as running out of memory on an 80G A100 GPU. The primary computational burden stems from the integration of local and deformable attention mechanisms, where local attention may not benefit resolving spatial misalignment when fusing defocus blur maps with blurry images.
4.6.2 Framework Components Ablation
To systematically evaluate the contributions of our reblurring-guided learning framework, we design ablation variants based on the UNet deblurring model, all trained on the SDD dataset: The effectiveness of the deblurring model involving pseudo defocus blur map for supervision has already been demonstrated in Table 1, where Ours with Eq. (2) outperforms Ours with Eq. (1). Our ablation studies are performed on the SDD dataset, and six variants are designed to analyze their contributions. Variant #1 represents a vanilla UNet trained using $\mathcal{L}_{1}$ loss. Variants #2 and #3 correspond to Ours with Eq. (1) by excluding the bi-directional optical flow deforming process and calibration mask, respectively. Variants #4, #5, and #6 aim to demonstrate the effectiveness of components within our reblurring module. Specifically, they variants of Ours with Eq. (1) by removing the reblurring module, isotropic kernels module $\mathcal{R}_{kpn}$ and weight prediction module $\mathcal{R}_{wpn}$ from, respectively.
@@ 4.6.2 Framework Components Ablation
#2 w/ Warp Operation 25.51/0.776/0.322
#3 w/o Cycle Deformation 24.80/0.758/0.347
#4 w/o Calibration Mask 25.78/0.780/0.309
#5 w/o reblurring module 25.58/0.777/0.311
#6 w/o $\mathcal{R}_{kpn}$ 25.68/0.778/0.311
#7 w/o $\mathcal{R}_{wpn}$ 25.74/0.778/0.332
Ours (Eq. (1)) 25.82/0.783/0.305
Ours (Eq. (2)) 26.20/0.796/0.298
The quantitative results are presented in Table 8. Observations from experiments #1, #2 and #3 demonstrate that the absence of the deformation process significantly diminishes network performance. Moreover, the inclusion of the calibration mask appears to moderately enhance performance, as depicted in experiment #4. Experiment #5 clearly indicates that the lack of reblurring module substantially impacts overall network performance. The results of experiments #6 and #7 validate the advantages of isotropic blur kernels prediction module $\mathcal{R}_{kpn}$ and weight prediction module $\mathcal{R}_{wpn}$ within reblurring module for the reblurring process.